# PEPSI Investigation, Retrieval, and Atlas of Numerous Giant Atmospheres (PIRANGA). III. Composition and winds in the atmosphere of TOI-1518 b

### v1

@@ 1 Introduction
Ultra-hot Jupiters (UHJs) are important objects for atmospheric study due to their close orbits to their host stars, large radii, and high temperatures. Transmission spectra enable the detection of elements and molecules in transiting exoplanets, searching for their absorption lines in the terminator, i.e. the transition region between the day and night side of the exoplanet. Transmission spectroscopy is most sensitive to atmospheres with a large scale height through which more stellar light is absorbed. UHJs have large scale heights due to their high equilibrium temperature and therefore provide high signal-to-noise ratio (SNR) data. The frequent transits of UHJs allow for flexibility in obtaining observations and enable the repeated measurements needed to verify atmospheric properties.
Among $\sim$19 UHJs with equilibrium temperatures $T_{eq}>2000$ K, radius $R>R_{J}$, and magnitude $V<10$, (e.g. West et al. 2016; Gaudi et al. 2017; Evans et al. 2017), TOI-1518 b is an interesting target for several reasons. First, the discovery paper found that the planet has a near-grazing transit with an impact parameter $\sim$0.9 (Cabot et al., 2021). They also detected Fe I and found evidence for Fe II in the transmission spectra and suggested that it may be a future target for emission spectroscopy to search for thermal inversions, where the atmospheric temperature increases with altitude. A search for a thermal inversion was later completed in Petz et al. (2024a), who used emission spectroscopy to confirm the existence of an inversion. The agent responsible for thermal inversions remains elusive. While molecules like TiO and VO are often suspected to be responsible (Hubeny et al., 2003; Fortney et al., 2008) for thermal inversions, they are very rarely detected (Nugroho et al., 2020; Prinoth et al., 2022; Johnson et al., 2023). They are likely photodissociated and therefore do not play an important role. There is increasing evidence that other metals, particularly Fe, which is universally detected in UHJ atmospheres, absorb in the ultraviolet (UV) and optical, offering a compelling alternative explanation to the inversion phenomena (Gandhi & Madhusudhan, 2019; Petz et al., 2024a).
@@
In this paper, we utilize high-resolution spectroscopic data from the PEPSI spectrograph at the Large Binocular Telescope (LBT) to (1) search for evidence of atomic and molecular species, (2) conduct time-resolved wind velocity measurements in the transmission spectra of TOI-1518 b, and (3) measure the transit parameters of TOI-1518b for comparison with the expected values and previously detected nodal precession. In §2, we describe our high-resolution spectroscopic observations. §3 describes our data reduction, telluric removal, and systematics removal. In §4, we describe our template spectra, cross-correlation functions, and Doppler shadow removal. Finally, §5 presents and discusses our detected species and measured wind velocities.
@@ 2 Observations
Data were obtained on 2022 September 26 UT using the Potsdam Echelle Polarimetric and Spectroscopic Instrument (PEPSI; Strassmeier et al. 2015) on the LBT. The LBT has dual 8.4m apertures with the light simultaneously fed from both primary mirrors to the spectrograph. Two different wavelength ranges were observed simultaneously: 4800-5441Å on the blue arm (CD3) and 6278-7419Å on the red arm (CD5). The resolving power of both is Rsimilar-$\sim$130,000 with a fiber diameter of 200 microns. A series of 50 exposures were taken over the course of the night, before, during, and after the transit of TOI-1518 b. Each observation had an exposure time of 250s with a median continuum SNR of 146 (blue arm) and 181 (red arm).
@@ 3 Data Reduction and Processing
We fit and removed telluric lines from the red arm using Molecfit (Kausch et al. 2015; Smette et al. 2015). Tellurics primarily affect the data from the red arm, while the blue arm is largely telluric-free (see top panel of Fig. 1).
@@ 4 Analysis
Table 1: Parameters of TOI-1518 from MISTTBORN. Priors are given by either a uniform ($\mathcal{U}$) or normal ($\mathcal{N}$) distribution. *We introduced a scale factor to MISTTBORN to account for the depth of the Doppler shadow compared to the input CCFs. The input CCFs are in SNR space. Dividing these by the peak SNR multiplied by some factor yields a Doppler shadow $\sim$1the depth of the normalized line profile. This scale factor also alters the reported $R_{P}/R_{\star}$, so the value we quote here is not directly comparable to $R_{P}/R_{\star}$ from the literature. ** $k=R_{P}/R_{\star}$**To avoid our model choosing an anomalously high v sin i, which would not be supported by independent constraints of the star’s rotation, we reduced the uncertainty on the prior by a factor of 100. This factor was determined by reducing the uncertainty by an order of magnitude until the model consistently adopted v sin i roughly equal to the literature value.
@@
We combine the results from both the blue and the red arms using a weighted sum to calculate the shifted CCF in the upper left panel of Fig. 3. The weight is the SNR of the spectrum squared multiplied by the total equivalent width of lines in the template spectrum. We iterate over a grid of radial velocity semi-amplitudes ($K_{P}$) to shift the CCFs into the planetary rest frame. The presence of Fe I is suggested by the red region centered on the crosshairs. The peak SNR is measured as the maximum SNR in a box within ±$\pm 20$km s-1 of these crosshairs.
Figure 2: (Top) Raw CCF for Fe I. Dark (black) regions are anti-correlated and light (white) regions are correlated. The blue lines show the ingress and egress of the transit, calculated using the transit duration determined by Duck et al. (in prep). The red dotted line shows the expected velocity of the atmospheric signal $V_{sig}=K_{P}\sin(2\pi\phi)+V_{sys}$, where $K_{P}=157$ km s-1 is the radial velocity semi-amplitude (Cabot et al. 2021), $\phi$ is the orbital phase, and $V_{sys}=-11.17$ km s-1 is the systemic velocity offset in the LBT PEPSI frame (Petz et al., 2024a). The dark shaded region is the Doppler shadow. It makes a V shape with the light shaded region, which is the observed atmospheric signal. (Middle) Model of the Doppler shadow from HORUS. This adopts transit parameters shown in Table 1 which were obtained via MCMC with MISTTBORN. (Bottom) Residual after subtracting the Doppler shadow.
@@ 5.1 Detecting Species
Fe I $5.4\times 10^{-4}$ 7.8 $-7.9\pm 1.0$
Fe II $5.4\times 10^{-3}$ 8.9 $-6.8\pm 1.0$
Cr I $7.4\times 10^{-7}$ 4.4 $-8.8\pm 2.1$
Ni I $2.8\times 10^{-7}$ 4.0 $-8.0\pm 1.0$
We define tentative detections as detections having peak SNR $>3\sigma$ and good detections as detections with peak SNR $>5\sigma$. Under this definition, we detected Fe I and Fe II using the combined blue and red arm data (Table 5.1). The corresponding shifted CCF plots are labeled in the upper panels of Fig. 3. We also tentatively detected Cr I and Ni I, with their shifted CCF plots shown in the lower panels of Fig. 3.
@@ 5.3 Time-Resolved Wind Velocities
We search for this effect by creating a phase-binned CCF plot. We first bin our observations based on orbital phase, choosing an initial bin size of 0.005. The aim is to have sufficient SNR in each bin to detect the time-resolved wind velocities, while also having a meaningful number of phase bins to observe changes over the course of the transit. The CCFs in each bin are summed using the CCF weights described in §4. We fit a Gaussian to each bin using the combined GP + MCMC analysis described in §5.2. We further check that the peak SNR > 3 or that the amplitude of these Gaussians show a SNR > 3. If they do not, we combine the data in this bin with the next bin forward in time and continue the analysis. This gives us a variable bin size which is increased during times of the transit where we have less available data or the data has lower SNR.
@@
divided by $\nu=6$ degrees of freedom (dof). Where there were missing values in one of the phase bins, we used linear interpolation to find the wind velocity and uncertainty from the nearest points. For the 6 data points of Fe I (Fe II), this gives a value of 1.6 (1.5). For 6 dof, a value $>$ 2.1 would reject the null hypothesis at a 95% confidence level (see, e.g., Pugh & Winslow 1966). This suggests that the wind velocities we measured are consistent with coming from the same underlying distribution.
Considering that we are missing some observations near ingress, our measurements of the time resolved wind velocities could be consistent anywhere from a similar interpretation as the winds in WASP-76 b (Ehrenreich et al., 2020) to a constant wind velocity. The SNR for Cr I and Ni I is not high enough to detect the time-resolved wind velocities in smaller phase bins.
@@ 6 Conclusions
The authors have been made aware that during the course of writing this paper, another paper by Simonnin et al. (2024) has appeared on arXiv with a complementary analysis of TOI-1518 b. They used MAROON-X to detect 14 species in the atmosphere of TOI-1518 b, also presenting 6 of these species, including Fe I and Fe II, with time-resolved wind velocities.
@@ Acknowledgements
MCJ and CL were supported by NASA Grant 80NSSC23K0692. This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.

### v2

@@ 1 Introduction
Ultra-hot Jupiters (UHJs) are important objects for atmospheric study due to their close orbits to their host stars, large radii, and high temperatures. Transmission spectroscopy is a powerful tool for this research, enabling the detection of atmospheric species by searching for their absorption lines in the terminator, i.e. the transition region between the day and night sides of the exoplanet. UHJs are particularly favorable targets for transmission spectroscopy because their high equilibrium temperatures directly lead to large atmospheric scale heights, which provides high signal-to-noise ratio (SNR) data. This high-quality data is crucial for exploring the extreme atmospheric properties of UHJs including large day-to-night temperature contrasts which drive strong day-to-night winds (e.g. Snellen et al. 2010; Ehrenreich et al. 2020; Pai Asnodkar et al. 2022), super-rotating equatorial jets (Showman et al. 2009; Showman & Polvani 2011), or high dayside temperatures which lead to extreme chemistry (Cooper & Showman 2006; Bell & Cowan 2018; Showman et al. 2020). It is also possible to distinguish the spatial locations of different atmospheric species, which can further aid in determining the properties and circulation patterns of the atmosphere. Detecting and cataloging additional atmospheric species is crucial to evaluating the chemistry of UHJ atmospheres. Furthermore, the frequent transits of UHJs offer flexibility in obtaining observations and enable the repeated measurements needed to verify these atmospheric properties.
[…]
@@
In this paper, we utilize high-resolution spectroscopic data from the PEPSI spectrograph at the Large Binocular Telescope (LBT) to (1) search for evidence of atomic and molecular species, (2) conduct time-resolved wind velocity measurements in the transmission spectra of TOI-1518 b, and (3) measure the transit parameters of TOI-1518b for preliminary comparison with the expected values and previously detected nodal precession. In §2, we describe our high-resolution spectroscopic observations. §3 describes our data reduction, telluric removal, and systematics removal. In §4, we describe our template spectra, cross-correlation functions, and Doppler shadow removal. Finally, §5 presents and discusses our detected species and measured wind velocities.
@@ 2 Observations
[…]
@@ 3 Data Reduction and Processing
We fit and removed telluric lines from the red arm using Molecfit (Kausch et al. 2015; Smette et al. 2015). The tellurics affecting the data from the red arm are shown in the top panel of Fig. 1. The blue arm is largely telluric-free, with only weak ozone absorption bands present. These are removed via the implementation of SYSREM below.
@@ 4 Analysis
Table 1: Parameters of TOI-1518 from MISTTBORN. Priors are given by either a uniform ($\mathcal{U}$) or normal ($\mathcal{N}$) distribution. *We introduced a scale factor to MISTTBORN to account for the depth of the Doppler shadow compared to the input CCFs. The input CCFs are in SNR space. Dividing these by the peak SNR multiplied by some factor yields a Doppler shadow $\sim$1% the depth of the normalized line profile. This scale factor also alters the reported $R_{P}/R_{\star}$, so the value we quote here is not directly comparable to $R_{P}/R_{\star}$ from the literature. ** $k=R_{P}/R_{\star}$ ***To avoid our model choosing an anomalously high v sin i, which would not be supported by independent constraints of the star’s rotation, we reduced the uncertainty on the prior by a factor of 100. This factor was determined by reducing the uncertainty by an order of magnitude until the model consistently adopted v sin i roughly equal to the literature value.
@@
We combine the results from both the blue and the red arms using a weighted sum to calculate the shifted CCF in the upper left panel of Fig. 3. The weight is the SNR of the spectrum squared multiplied by the total equivalent width of lines in the template spectrum. This weighting scheme accounts for both observational noise from our spectra, as well as the strength of the signal expected for the red arm and the blue arm based on the lines in our templates. We iterate over a grid of radial velocity semi-amplitudes ($K_{P}$) to shift the CCFs into the planetary rest frame. The presence of Fe I is suggested by the red region centered on the crosshairs. The peak SNR is measured as the maximum SNR in a box within $\pm 20$km s-1 of these crosshairs.
Figure 2: (Top) Raw CCF for Fe I. Dark (black) regions are anti-correlated and light (white) regions are correlated. The blue lines show the ingress and egress of the transit, calculated using the transit duration determined by Duck et al. (in prep). The red dotted line shows the expected velocity of the atmospheric signal $V_{sig}=K_{P}\sin(2\pi\phi)+V_{sys}$, where $K_{P}=157$ km s-1 is the radial velocity semi-amplitude (Cabot et al. 2021), $\phi$ is the orbital phase, and $V_{sys}=-11.17\pm 0.035$ km s-1 is the systemic velocity offset in the LBT PEPSI frame (Petz et al., 2024a). The dark shaded region is the Doppler shadow. It makes a V shape with the light shaded region, which is the observed atmospheric signal. (Middle) Model of the Doppler shadow from HORUS. This adopts transit parameters shown in Table 1 which were obtained via MCMC with MISTTBORN. (Bottom) Residual after subtracting the Doppler shadow.
@@ 5.1 Detecting Species
Fe I $5.4\times 10^{-4}$ 7.8 $7.9\pm 1.0$
Fe II $5.4\times 10^{-3}$ 8.9 $6.8\pm 1.0$
Cr I $7.4\times 10^{-7}$ 4.4 $8.8\pm 2.1$
Ni I $2.8\times 10^{-7}$ 4.0 $8.0\pm 1.0$
We define tentative detections as detections having peak SNR $3\sigma$ and good detections as detections with peak SNR $5\sigma$. Under this definition, we detected Fe I and Fe II using the combined blue and red arm data (Table 5.1). The corresponding shifted CCF plots are labeled in the upper panels of Fig. 3. We also tentatively detected Cr I and Ni I, with their shifted CCF plots shown in the lower panels of Fig. 3.
@@ 5.3 Time-Resolved Wind Velocities
We search for this effect by creating a phase-binned CCF plot. We first bin our observations based on orbital phase, choosing an initial bin size of 0.005. The aim is to have sufficient SNR in each bin to detect the time-resolved wind velocities, while also having a meaningful number of phase bins to observe changes over the course of the transit. The CCFs in each bin are summed using the CCF weights described in §4. We fit a Gaussian to each bin using the combined GP + MCMC analysis described in §5.2. We further check that the peak SNR > 3 or that the amplitude of these Gaussians show a SNR > 3. If they do not, we combine the data in this bin with the next bin forward in time and continue the analysis. This gives us a variable bin size which is increased during times of the transit where we have less available data or the data has lower SNR. We run tests in binning the data with forward and reverse time and confirm that the binning strategy does not affect the conclusions in this paper.
@@
divided by $\nu=6$ degrees of freedom (dof). Where there were missing values in one of the phase bins, we used linear interpolation to find the wind velocity and uncertainty from the nearest points. For the 6 data points of Fe I (Fe II), this gives a value of 1.3 (1.6). For 6 dof, a value >> 2.1 would reject the null hypothesis at a 95% confidence level (see, e.g., Pugh & Winslow 1966). This suggests that the wind velocities we measured are consistent with coming from the same underlying distribution.
Considering that we are missing some observations near ingress, our measurements of the time resolved wind velocities could be consistent anywhere from a similar interpretation as the winds in WASP-76 b (Ehrenreich et al., 2020) to a constant wind velocity. We perform a statistical test to check if the data prefers (1) a vertical constant line model or (2) a toy model with a linear change in velocity followed by a constant velocity. The second model would be consistent with the physical interpretation from Ehrenreich et al. (2020). For Fe I, the reduced $\chi^{2}$ values are 1.6 and 2.3 for Model #1 and #2. For Fe II, the reduced $\chi^{2}$ values are 1.0 and 0.7 for Model #1 and #2. Based on these values and the uncertainty of error bars, we cannot draw a strong conclusion as to whether our data prefers a vertical constant model or a linear+constant toy model. The SNR for Cr I and Ni I is not high enough to detect the time-resolved wind velocities in smaller phase bins.
@@ 6 Conclusions
The authors have been made aware that during the course of writing this paper, another paper by Simonnin et al. (2024) has appeared on arXiv with a complementary analysis of TOI-1518 b. They used MAROON-X to detect 14 species in the atmosphere of TOI-1518 b including Fe I, Fe II, and Cr I as identified in this paper, but not including Ni I. Their detection of Ti I is also of interest as an alternative to detecting TiO, which has been elusive to date despite being a candidate to cause temperature inversions. Simonnin et al. (2024) also present 6 of these species (Fe I, Fe II, Ca I, Ca II, Na I, and Mg I) with time-resolved wind velocities. The velocity patterns of Fe I and Fe II are qualitatively consistent between the two papers, with a constant or slightly increasing blueshift over part of the transit. However, our missing or poor quality observations during ingress preclude a full comparison.
@@ Acknowledgements
MCJ, JW, and CL were supported by NASA Grant 80NSSC23K0692. This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.
The LBT is an international collaboration among institutions in the United States, Italy and Germany. LBT Corporation Members are: The Ohio State University, representing OSU, University of Notre Dame, University of Minnesota and University of Virginia; LBT Beteiligungsgesellschaft, Germany, representing the Max-Planck Society, The Leibniz Institute for Astrophysics Potsdam, and Heidelberg University; The University of Arizona on behalf of the Arizona Board of Regents; and the Istituto Nazionale di Astrofisica, Italy. Observations have benefited from the use of ALTA Center (alta.arcetri.inaf.it) forecasts performed with the Astro-Meso-Nh model. Initialization data of the ALTA automatic forecast system come from the General Circulation Model (HRES) of the European Centre for Medium Range Weather Forecasts.